# Bayesian estimation of cross-coupling and reflection systematics in 21cm array visibility data

### v1

@@ 4 Modelling
This section details the modelling of the foregrounds and EoR signals, as well as the setup of the Bayesian model estimation problem.
@@ 4.1 Foreground Model
In order to avoid having to model observed foregrounds to high precision, HERA uses a foreground-avoidance technique in its observations (Kerrigan et al., 2018; Morales et al., 2018), which aims to retain only the Fourier modes outside of the foreground wedge, i.e. regions where the foregrounds are less dominant. This foreground wedge is typically defined in 2D Fourier space – the plane formed by $k_{\parallel}$ and $k_{\bot}$, where the former is dependent on the spectral resolution and bandwidth of the array, and the latter on the baseline distribution. However, in this work, an assumed foreground model is required, both to construct the visibilities, and copies of which are added back in at other delays due to the systematics.
Figure 3: An example of ten cross-correlation foreground eigenvectors for a 14.7 m east-west baseline, with the top being the real components, and the bottom the imaginary. The autocorrelation foregrounds exhibit similar eigenvector characteristics, albeit with only the real components being non-zero.
Figure 4: The first fifty cross-correlation foreground eigenvalues found through the eigenvalue decomposition.
@@
The simplified model used here is constructed from simulated HERA Validation foregrounds, which consist of the GaLactic and Extragalactic All-sky Murchison Widefield Array (GLEAM) point-source catalogue (Hurley-Walker et al., 2017), with additional bright sources added, and a diffuse component based on the extended Global Sky Model (eGSM, Zheng et al., 2017; Kim et al., 2018). It was sufficient in our case (focusing on the shortest baselines) to construct the foreground model from the diffuse component only. Modelling the point-source components as well would likely require a number of higher-order modes, increasing the already large number of model parameters.
@@
Fig. 3 plots the first ten eigenvectors for the cross-coupling foreground model. Bandpass gains were applied to the foregrounds prior to the eigendecomposition (but were not calibrated for). The strongest modes model the most prominent features in the foregrounds, which in this case is the smoothest structures across frequency. Fig. 4 shows the first fifty eigenvalues from the decomposition. The behaviour is as expected, with there being a rapid decrease in eigenvalue amplitudes towards higher modes. Fig. 5 shows the covariance and correlation matrices for the simulated diffuse foregrounds from which our eigenmodes are constructed.
The foreground model used here does not include any EoR contributions. The EoR is modelled separately, and is discussed in Section 4.2.
In order to effectively model the autocorrelation foregrounds down to the EoR level (after which the EoR power dominates), thirty eigenvectors were needed. For cross-correlations, one hundred eigenvectors were needed. In terms of the difference of seventy eigenvectors between the autocorrelations and cross-correlations, this is likely due to cross-correlations containing more spectral structure than the autocorrelations. The autocorrelations are symmetric around $\tau=0$, while the cross-correlations are not, so the latter requires additional modes in order to accurately recover the foreground power down to the EoR level at both positive and negative delays.
In order to determine the number of required eigenvectors, a basis projection between the eigenvectors and an example true foreground was carried out. This projection is simply $\lambda_{\rm proj}={x_{\rm{FG}}}\cdot{V_{\rm{FG,true}}}$, where $\lambda_{\rm proj}$ are the eigenvalues following projection, found by the dot product of our eigenvectors ($x_{\rm{FG}}$) and a true foreground ($V_{\rm{FG,true}}$). Having too few eigenvectors resulted in poor recovery of the true power spectrum, as well as excess power in regions dominated by systematics and the EoR. This would result in compromised modelling of these two components, and could limit the recovery of the underlying 21-cm signal.
Fig. 6 shows the reconstructed foreground power spectrum using these eigenvalues when using either 50 or 100 of the highest amplitude modes for a cross-correlation foreground. The EoR power spectrum is included for comparison. The red dot-dashed line shows that with 50 eigenvectors the main foreground peak can be recovered, but not the lower-amplitude, higher-frequency features. This is problematic as the recovered foreground diverges from the true foreground at delays where the EoR signal begins to dominate. Even though the recovered foreground is sub-dominant at $\tau\gtrsim 750\,\rm{ns}$, this could lead to an overestimation of the power contribution from the foregrounds, with an associated underestimation of the EoR and systematics contributions.
Figure 6: Examples of the true foreground recovery achievable following a projection of our foreground eigenvectors onto a true foreground (in a 14.7 m east-west cross-correlation power spectrum). The black solid line is an example true foreground, and the grey dashed line an EoR signal which is included to demonstrate where it begins to dominate in power. The red dot-dashed line shows the recovery when using 50 foreground eigenvectors, and the blue dash-dot-dotted line 100 eigenvectors.
Increasing the number of eigenvectors to 100 results in a better recovery of the true foregrounds (blue dash-dot-dotted line). While not perfect, any deviations are well-enough below the EoR’s power level, and would have minimal effect on recovery of the 21-cm signal. While there is certainly a compromise to be found between 50 and 100 modes which would provide acceptable modelling of the foregrounds, a total of 100 eigenvectors was settled on as there was good recovery, and an $\mathcal{O}$(10) increase to the number of modelling parameters would negligibly affect the Bayesian model performance in this context.
[…]
@@ 4.2 EoR Model
The modelling here also requires at least a simplistic EoR model to account for the approximately constant power in regions not dominated by foregrounds (i.e. $\tau\gtrsim 750\,\rm{ns}$ in Fig. 6), so it is not sufficient to only model the foregrounds and systematics despite the fact that copies of the foreground dominate said systematic effects. The EoR model used here is formed from the full-sky, full-bandwidth, and physically motivated mock-EoR simulated in the HERA Validation simulations (Aguirre et al., 2022), where it is modelled as a Gaussian random brightness temperature field.
Figure 7: Left: The autocorrelation power spectrum of the noise estimate (grey dashed line) for a single time integration in comparison to the true EoR (red dot-dashed line) and true foreground (blue solid line) power spectra across the delay range of interest. Right: A comparison with the noise estimate in the cross-correlation regime. The noise estimate is derived from Eq. 14.
@@
For a particular baseline and polarisation, an example EoR visibility is Fourier transformed, and $200$ Fourier modes are taken. This number was chosen as it provides good coverage of the delay range $(-2000,+2000)$ ns, which contains the entirety of the foreground peak and all of the systematics considered in this work. Fewer modes would result in gaps in delay space. Our model simply fits each of the 200 Fourier modes individually with 200 amplitude parameters. In other words, we do not fit a fixed guess of the EoR visibility which has a predetermined structure with a single amplitude parameter. This is discussed further in Section 4.4.
@@ 4.3 The No-U-Turn Sampler
Each chain in the sampler is then tuned for 5000 steps, and sampled for a total of 15000 steps with a target acceptance rate of 0.9 (increased from the default 0.8). When sampling, the step size used in the model is adjusted in order to achieve the given target acceptance rate. A larger target acceptance can help with sampling of complex posteriors (e.g. those which deviate from a Gaussian in shape), but can also result in slower sampling (Betancourt, 2017).
@@
$\boldsymbol{\sigma}_{ij}=\sqrt{\frac{V_{ii}V_{jj}}{\Delta t\Delta\nu}},$ (14)
where $\Delta\nu=97.7$ kHz is the frequency resolution of the simulated data, with an integration time of $\Delta t=10.73$ seconds. $V_{ii}$ and $V_{jj}$ are the autocorrelations visibilities for the two antennas in a particular baseline when analysing a cross-correlation.
@@ 4.4 Priors
In this section we outline our prior choice methodology. These are our best estimates for the systematics parameters before the model is sampled. When generating the mock data, a level of randomisation was introduced in all systematics parameters. As such, it was assumed that we did not have perfect knowledge of the systematics’ characteristics. Coupled with the noise, estimates were chosen to take this randomisation into account. For example, if a power spectrum was visually inspected, and a reflection systematic was estimated to be located at a delay of 200 ns, the prior was set to be a normal distribution centred on this delay, with a standard deviation which takes into account the uncertainty of its true delay, the noise, and perhaps any confusion between closely neighbouring systematics. The sampler then evaluates these priors against the mock data in order to form the posterior distributions, which will, ideally, be centred on the true delay.
@@
Foreground eigenvector amplitudes (100) $10^{-1}-10^{-6}$ $5\%$
EoR Fourier mode amplitudes (200) $1$ $500\%$
@@
Table 1: A description of all the priors used in the model, as well as how many of each parameter there are, and their units. All priors are normal distributions. For brevity, only the ranges and orders of magnitudes of some parameters are given. This is to provide a broad idea of how the prior widths compare to the means, and is due to fact that an exhaustive list would likely not be informative. Cross-coupling foregrounds needed significantly more eigenvectors than the autocorrelations, likely to more accurately model the additional spectral structure. The foreground amplitude prior means were derived from a vector projection of the model eigenvectors. The EoR amplitude prior means are all simply $1$, as the Fourier modes derived from an example EoR visibility already showed good agreement with the EoR component in the mock data. Systematic amplitudes are measured relative to the foreground peak, and delays are in units of seconds. All phase priors are $\mathcal{N}(\mu=0,\sigma=1)$, which covers the $[-\pi,\pi)$ range. Subreflection amplitude prior widths are wider than other systematics as it is not immediately evident through visual inspection how much of the amplitude at a given delay is from the subreflection in question, neighbouring subreflections, or the foreground and EoR components. Note that the prior means are not necessarily identical to the parameters used to generate the systematics. When generated, randomisation is introduced, so the priors can be best initial guesses with this in mind.
@@
Furthermore, fair knowledge of the foregrounds were assumed (in terms of prior specificity). The amplitudes of the prior means for the autocorrelation foregrounds ranged from $\mathcal{O}(10^{-3})-\mathcal{O}(10^{2})$. Normal prior standard deviation widths of $5\%$ the amplitudes were chosen. Cross-correlations are lower in power, with the absolute values of these amplitudes ranging from $\mathcal{O}(10^{-6})-\mathcal{O}(10^{-1})$. The prior standard deviations were reduced in order to maintain a relative width of $5\%$.
Since the EoR model was constructed from the Fourier transform of a true EoR visibility, the amplitudes of the modes immediately provided good agreement, and so did not require significant adjustments. As such, prior means of $1$ were used for the amplitudes of the EoR Fourier modes. It is, however, less reasonable to assume good knowledge of the EoR signal’s properties, given the difficulty associated with its observation. As such, prior widths $500\%$ of the amplitudes were chosen.
@@
Lastly, for the systematics in question, our model consists of 66 free parameters for the reflection systematics, and 30 for cross-coupling. While the implementation of Kern et al. (2019)’s method differs for newer Phase I observational data, we compare to the results when applied to early Phase I data, wherein the reflections were fitted with 28 terms (84 free parameters), and the cross-coupling model was formed with the first 30 SVD modes (Kern et al., 2020a; The HERA Collaboration et al., 2023). The discrepancy between the number of reflection free parameters in comparison to our model is likely due to the fact that we follow the setup of Aguirre et al. (2022), which is not a perfect simulation of the early Phase I observational data. Kern et al. (2020a) likely just saw an increased number of subreflections, or required additional terms in order to achieve sufficient subtraction.
@@ 4.5 Systematics Subtraction
$V_{apbq,t^{\prime}}^{\rm{recovered}}=(\tilde{g}_{b})^{-1}(\tilde{g}_{a})^{-1}V_{apbq,t}^{\rm{refl}}.$ (16)
This is done with all 30,000 samples from the model. This produces 30,000 recovered visibilities, from which the 95th percentile can be formed.
For cross-coupling, two models are formed from the results: the first containing only model foregrounds and EoR, and the second containing model foregrounds, EoR, and cross-coupling. More specifically, using the foreground eigenvector amplitudes ($\lambda^{\rm{Model}}$), and EoR Fourier mode amplitudes ($A^{\rm{Model}}$) from the sampling results, a systematics-free visibility is formed. With the foreground eigenvectors ($\textbf{x}_{\rm{FG}}$) from Section 4.1, and the EoR Fourier modes ($f_{\rm{EoR}}$) from Section 4.2, this is formed with
$V_{\rm{FG+EoR}}^{\rm{Model}}=\lambda^{\rm{Model}}\textbf{x}_{\rm{FG}}+A^{\rm{Model}}f_{\rm{EoR}}.$ (17)
@@
$V^{\rm{Model}}_{\rm{Corrupt}}=\lambda^{\rm{Model}}\textbf{x}_{\rm{FG}}+A^{\rm{Model}}f_{\rm{EoR}}+V_{\rm{CC}}^{\rm{Model}}$ (18)
Taking the difference of Eqs. 17 and 18 produces a cross-coupling-only visibility,
$V_{\rm{CC\;Only}}^{\rm{Model}}=V^{\rm{Model}}_{\rm{Corrupt}}-V_{\rm{FG+EoR}}^{\rm{Model}}.$ (19)

### v2

@@ 4 Modelling
This section details the modelling of the foregrounds and EoR signals, as well as the setup of the Bayesian model estimation problem and model [^2] https://github.com/GeoffMurphy/HMCSystematicsSampler/.
@@ 4.1 Foreground Model
In order to avoid having to model observed foregrounds to high precision, HERA uses a foreground-avoidance technique in its observations (Kerrigan et al., 2018; Morales et al., 2018), which aims to retain only the Fourier modes outside of the foreground wedge, i.e. regions where the foregrounds are less dominant. This foreground wedge is typically defined in 2D Fourier space – the plane formed by $k_{\parallel}$ and $k_{\bot}$, where the former is dependent on the spectral resolution and bandwidth of the array, and the latter on the baseline distribution. However, in this work, an assumed foreground model is required, both to construct the visibilities, and to simulate the systematics which produce copies of these foregrounds at higher delays.
Figure 3: An example of ten cross-correlation foreground eigenvectors for a 14.7 m east-west baseline, with the top being the real components, and the bottom the imaginary. These ten modes are the eigenvectors corresponding to the largest eigenvalues in the decomposition, with “Mode 0” corresponding to eigenvalue 0 in Fig. 4, “Mode 1” to eigenvalue 1, etc. Furthermore, they have the same units as the foregrounds in frequency space. The autocorrelation foregrounds exhibit similar eigenvector characteristics, albeit with only the real components being non-zero.
Figure 4: The first fifty cross-correlation foreground eigenvalues found through the eigenvalue decomposition. These are unitless, as they are factors multiplied with their corresponding foreground eigenvector, which subsequently forms a model foreground in units of Janskys.
@@
The simplified model used here is constructed from simulated HERA Validation foregrounds, which consist of the GaLactic and Extragalactic All-sky Murchison Widefield Array (GLEAM) point-source catalogue (Hurley-Walker et al., 2017), with additional bright sources added, and a diffuse component based on the extended Global Sky Model (eGSM, Zheng et al., 2017; Kim et al., 2018). It was sufficient in our case (focusing on the shortest baselines) to construct the foreground model from the diffuse component only. Modelling the point-source components as well would likely require a number of higher-order modes, increasing the number of model parameters.
@@
Fig. 3 plots the first ten eigenvectors for the cross-coupling foreground model. Bandpass gains were applied to the foregrounds prior to the eigendecomposition (but were not calibrated for). The strongest modes model the most prominent features in the foregrounds, which in this case is the smoothest structures across frequency. Fig. 4 shows the first fifty eigenvalues from the decomposition. The behaviour is as expected, with there being a decrease in eigenvalue amplitudes towards higher modes. Fig. 5 shows the covariance and correlation matrices for the simulated diffuse foregrounds from which our eigenmodes are constructed. Eight eigenvectors were required to sufficiently recover the foregrounds, for both the autocorrelations and cross-correlations.
@@ 4.2 EoR Model
The model was also tested with and without an included signal component. Initially, it was thought that the approximately constant power across delay needed to be accounted for (see Fig. 6), but it was found that the results were comparable between both cases. This is due to the power of the systematics, which are much higher than the signal, so the lack of a signal component does not adversely affect modelling to any significant degree. The EoR model tested here is formed from the full-sky, full-bandwidth, and physically motivated mock-EoR simulated in the HERA Validation simulations (Aguirre et al., 2022), where it is modelled as a Gaussian random brightness temperature field.
@@
For a particular baseline and polarisation, an example EoR visibility is Fourier transformed, and $200$ complex Fourier modes are uniformly sampled from the delay range $(-2000,+2000)$ ns, which contains the entirety of the foreground peak and all of the systematics considered in this work. Fewer modes would result in gaps in delay space. Our model simply fits each of the 200 Fourier modes individually with 200 amplitude parameters. This is discussed further in the context of priors in Section 4.6.
It was found that in realistic, noisy data, the recovered power spectra from the signal-free model was essentially equivalent to that found with a model containing a signal.
While the inclusion of a signal model slightly improved modelling in essentially noise-free cases (namely in the form of narrower uncertainty estimates in some cases), as it is implemented here with 200 Fourier modes, the run-times became prohibitive. This is due to a combination of the low noise, as well as the number of added parameters. This improvement also presented itself as better recovery of the underlying power spectra, but for reflection modelling this was a negligible gain in the very-low noise regime (i.e. it was comparable to the result of Fig. 10). For cross-coupling mitigation with very-low noise, the recovery improved by approximately half an order of magnitude with the inclusion of a signal model. Overall, the benefits of the inclusion of this component were outweighed by the performance loss in negligible-noise scenarios, and its inclusion was not necessary in high noise regimes.
Nevertheless, realistic-noise cases are more relevant. Here, the inclusion of a signal component provides comparable results to a signal-free model. As such, we opt to exclude this component in our analysis, which also allows for faster sampling speed.
Figure 6: Left: The autocorrelation power spectrum of the noise estimate (grey dashed line) for a single time integration in comparison to the true EoR (red dot-dashed line) and true foreground (blue solid line) power spectra across the delay range of interest. Right: A comparison with the noise estimate in the cross-correlation regime, in this case for a 14.7 m east-west baseline. The noise estimate is derived from Eq. 15.
4.3 The Systematics Model
Combining the foreground and systematics terms in Eqs. 7 and 11, respectively, we get the full systematics equation for our model:
$V^{\rm{corrupt}}_{apbq,t}=\sum\limits_{m=1}^{\rm{M}_{\rm{fg}}}\lambda_{abm}x_{abm}+\begin{cases}\tilde{g}_{a}\tilde{g}_{b}^{*}V_{apbq,t^{\prime}}&a=b\\ \tilde{g}_{a}\tilde{g}_{b}^{*}V_{apbq,t^{\prime}}+V_{apbq,t}^{\rm{cc}}&a\neq b\end{cases},$ (13)
where
$V_{apbq,t}^{\rm{cc}}=\tilde{g}_{a}\tilde{g}_{b}^{*}V_{apap,t^{\prime}}\left(\sum\limits_{j}^{\rm{N}_{CC}}{\rm{A}}_{apbq}^{d,j}\exp(i2\pi\nu\tau_{apbq}^{d,j}+i\phi_{apbq}^{d,j})\right)_{d\ni t}.$ (14)
Here, $x_{abm}$ are the foreground eigenvectors, and $\lambda_{abm}$ the corresponding free parameters for the foregrounds measured by antennas $a$ and $b$, with there being a total of $\rm{M}_{\rm{fg}}=8$ such modes in our model. As described in Section 2.1, $\tilde{g}_{a}$ and $\tilde{g}_{b}$ are our reflection gain terms (written in full in Eq. 6), which includes both high-amplitude cable reflections and subreflections (i.e. $\rm{M}=22$ in Eq. 6 for the fiducial model).
${\rm{A}}_{apbq}^{d,j}$ are the cross-coupling relative amplitudes described in detail in Section 2.3. Along with the delay, $\tau_{apbq}^{d,j}$, and phase, $\phi_{apbq}^{d,j}$, parameters, a total of $\rm{N_{\rm{CC}}}=10$ such features are modelled, where only the contribution from one of the two antennas are considered in this work, i.e. the coupling features present at positive delays. We also assume the coupling systematics undergo cable reflections, which accounts for the $\overline{g}_{a}\overline{g}_{b}^{*}$ factors in Eq. 14. As in the previous paragraph, there are 22 such reflection/gain terms. These are, however, typically very weak, with relative amplitudes of $10^{-7}-10^{-10}$, and so are negligible, but are still present in our model as result of the order in which systematics are added, i.e. the model adds coupling systematics first, followed by reflections, so reflected coupling is a natural byproduct of this.
Since we only model positive delays, we neglect the third term for the $a\neq b$ case. This would consist of the visibility from the second antenna, $V^{\rm{uncal}}_{bqbq,t^{\prime}}$, with distinct amplitudes, delays and phases for the coupling model. There would also be a leading factor of $\tilde{g}_{b}\tilde{g}_{a}^{*}$ denoting the cable reflection of the coupling-affected visibility.
4.4 Noise
$\boldsymbol{\sigma}_{ij}=\sqrt{\frac{V_{ii}V_{jj}}{\Delta t\Delta\nu}},$ (15)
where $\Delta\nu=97.7$ kHz is the frequency resolution of the simulated data, with an integration time of $\Delta t=10.73$ seconds. $V_{ii}$ and $V_{jj}$ are the autocorrelation visibilities for the two antennas in a particular baseline when analysing a cross-correlation.
@@ 4.5 The No-U-Turn Sampler
Each chain in the sampler is then tuned for 5,000 steps, and sampled for a total of 10,000 steps with a target acceptance rate of 0.8. When sampling, the step size used in the model is adjusted in order to achieve the given target acceptance rate. A larger target acceptance can help with sampling of complex posteriors (e.g. those which deviate from a Gaussian in shape), but can also result in slower sampling (Betancourt, 2017).
@@ 4.6 Priors
In this section we outline our prior choice methodology. These are our best estimates for the model parameters before sampling. When generating the mock data, a level of randomisation was introduced in all systematics parameters. As such, it was assumed that we did not have perfect knowledge of the systematics’ characteristics. Coupled with the noise, estimates were chosen to take this randomisation into account. For example, if a power spectrum was visually inspected, and a reflection systematic was estimated to be located at a delay of 200 ns, the prior was set to be a normal distribution centred on this delay, with a standard deviation which takes into account the uncertainty of its true delay, the noise, and perhaps any confusion between closely neighbouring systematics. The sampler then evaluates these priors against the mock data in order to form the posterior distributions, which will, ideally, be centred on the true delay.
@@
Foreground eigenvector amplitudes (8) $10^{-1}-10^{-3}$ $5\%$
@@
Table 1: A description of all the priors used in the model, as well as how many of each parameter there are, and their units. All priors are normal distributions. For brevity, only the ranges and orders of magnitudes of some parameters are given. This is to provide a broad idea of how the prior widths compare to the means, and is due to fact that an exhaustive list would likely not be informative. The foreground amplitude prior means were derived from a vector projection of the model eigenvectors. Systematic amplitudes are measured relative to the foreground peak, and delays are in units of seconds. All phase priors are $\mathcal{N}(\mu=0,\sigma=1)$, which covers the $[-\pi,\pi)$ range. Subreflection amplitude prior widths are wider than other systematics as it is not immediately evident through visual inspection how much of the amplitude at a given delay is from the subreflection in question, neighbouring subreflections, or the foreground and EoR components. Note that the prior means are not necessarily identical to the parameters used to generate the systematics. When generated, randomisation is introduced, so the priors can be best initial guesses with this in mind.
Table 1 describes the priors used in the sampler. When choosing priors, the amplitude and delay of a particular systematic can be fairly easily estimated. By Fourier transforming an observational visibility to delay space, the amplitude of the systematic peak can be compared to the foreground peak, and a relative amplitude can be inferred. Similar can be said of the delay of the systematic, where one can simply estimate the delay of the systematic peak visually, or from knowledge of the array’s geometry. In observational data, this is naturally complicated by noise, data cuts, etc.
@@
Furthermore, fair knowledge of the foregrounds were assumed (in terms of prior specificity). The amplitudes of the prior means for the autocorrelation foregrounds ranged from $\mathcal{O}(10^{2})-\mathcal{O}(10^{-3})$. Normal prior standard deviation widths of $5\%$ the amplitudes were chosen. Cross-correlations are lower in power, with the absolute values of these amplitudes ranging from $\mathcal{O}(10^{-1})-\mathcal{O}(10^{-3})$. The prior standard deviations were reduced in order to maintain a relative width of $5\%$.
@@
Lastly, for the systematics in question, our model consists of 66 free parameters for the reflection systematics, and 30 for cross-coupling. When implementing the reflection calibrator of Kern et al. (2019), one supplies the upper and lower delay bounds of each individual reflection peak, along with the number of iterations desired for its suppression. The coupling calibration requires the upper and lower bound of the entire coupling shelf, and the number of SVD modes to be used. For the mock data in this work, we opted for 15 SVD modes.
@@ 4.7 Systematics Subtraction
$V_{apbq,t^{\prime}}^{\rm{recovered}}=(\tilde{g}_{b}^{*})^{-1}(\tilde{g}_{a})^{-1}V_{apbq,t}^{\rm{refl}}.$ (18)
This is done with all 10,000 samples from the model. This produces 10,000 recovered visibilities, from which the 95th percentile can be formed.
For cross-coupling, two models are formed from the results: the first containing only model foregrounds, and the second containing model foregrounds and cross-coupling. More specifically, using the foreground eigenvector amplitudes ($\lambda^{\rm{Model}}$) from the sampling results, a systematics-free visibility is formed. With the foreground eigenvectors ($\textbf{x}_{\rm{FG}}$) from Section 4.1, this is formed with
$V_{\rm{FG}}^{\rm{Model}}=\lambda^{\rm{Model}}\textbf{x}_{\rm{FG}}.$ (19)
@@
$V^{\rm{Model}}_{\rm{Corrupt}}=\lambda^{\rm{Model}}\textbf{x}_{\rm{FG}}+V_{\rm{CC}}^{\rm{Model}}$ (20)
Taking the difference of Eqs. 19 and 20 produces a cross-coupling-only visibility,
$V_{\rm{CC\;Only}}^{\rm{Model}}=V^{\rm{Model}}_{\rm{Corrupt}}-V_{\rm{FG}}^{\rm{Model}}.$ (21)